Mardi recieved an inheritance of $70,000. She invested part at 10% and deposited the remainder in tax-free at 7%. Her total annual income from the investmentswas $6400. Find the amount invested at 10%.
Answer by richwmiller(17219)

x+y=70000,
0.1*x+0.07*y=6400
x=70000-y
0.1*(70000-y)+0.07*y=6400
7000-0.1y+0.07*y=6400
-0.03*y=-600
y=20000 at 7%
x=70000-y
x=50000 at 10%
check
0.1*50000+0.07*20000=6400
5000+1400=6400
6400=6400
ok